depth of 2 digit 10 to 99


1
Expert's answer
2021-02-24T12:51:57-0500

Number of digits from 10 to 99 is 90

and sum of digits from 10 to 99 is=

"S_{n}= \dfrac{n}{2}(2a + (n-1)d)"

Here,

n= 90

a= 10

d= 1

So, "S_{90}= \dfrac{90}{2}((2\times 10)+ 89\times1)"

"S_{90}= 45(20+ 89)= 45\times109"

"= 4905"

Hence, Sum of digits from 10 to 99 is 4905